# Evolution and compression in LLMs: On the emergence of human-aligned categorization

### v3

@@ 1 Introduction
As large language models (LLMs) become increasingly popular in everyday use, it is crucial to understand how their learning biases and representational capacities align with our own. Here, we investigate this by focusing on a key aspect of human intelligence: the ability to organize information into semantic categories (Croft2002Typology; Rosch2002Principlesa; Koch200885; Boster2005Categoriesa; Majid2015Comparing; Malt2013How). This phenomenon presents two major challenges for AI. First, systems of semantic categories (semantic systems, for short) exhibit both universal patterns and cross-language differences (Berlin1969Basic; Berlin1992Ethnobiological; Croft2002Typology), which LLMs must navigate. Second, LLMs are not grounded in the rich physical and social environment that humans are (Rosch1975Cognitive; Labov1973Boundariesa), and it is unclear how these differences affect their ability to learn human-aligned semantic categories. Therefore, in order to understand whether LLMs can efficiently communicate with people and adapt to changing environments and communicative needs in a human-like manner, it is crucial to study whether LLMs are capable of structuring meaning according to the same principles that guide humans.
@@
To address this open gap in our understanding of LLMs, we propose a novel theoretical and cognitively-motivated framework for studying semantic systems in LLMs. We build on the framework of Zaslavsky2018Efficient, which argues that languages efficiently compress meanings into words by optimizing the Information Bottleneck (IB) principle (Tishby1999Information), instantiated as a tradeoff between the informational complexity and communicative accuracy of the lexicon. This framework has broad empirical support across human languages (Zaslavsky2018Efficient; Zaslavsky2019Semantic; Zaslavsky2021Lets; Mollica2021Forms; Zaslavsky2022Evolution). Furthermore, Imel2025Iterated recently showed that a drive for IB-efficiency may be present in the individual inductive biases of human learners. While LLMs are trained on vast amounts of human language data encoded as text, they are not trained with respect to the IB objective. This raises the question: are LLMs capable of developing IB-efficient human-aligned semantic systems?
On the one hand, if LLMs fail to develop efficient semantic systems, it highlights a key misalignment in their capacity to learn human-like language. Furthermore, human semantic efficiency is fundamentally constrained by contingent, often perceptually grounded representations of meaning. Therefore, testing LLMs for efficiency also provides a test of whether they capture the specific representations humans use for categorization.
@@
We address this open question with an in-depth analysis in the domain of color—a key test case for categorization theories in cognitive science with rarely available human data as well as practical implication for human-LLM interactions (see Section 2.1)—and replicate with LLMs two influential human behavioral experiments (Figure 1). First, we conduct an English color naming experiment (analogous to the human experiment of Lindsey2014Color), designed to assess the efficiency and human-alignment of the color naming systems of LLMs. Second, we conduct an iterated learning experiment of pseudo color-naming systems (analogous to the human experiment of Xu2013Cultural), designed to reveal implicit inductive learning biases of LLMs by simulating a process of cultural transmission (see Section 2.3). For the latter, we extend Zhu2024Eliciting’s iterated in-context learning (I-ICL) paradigm to iterated in-context language learning (IICLL).
@@
Our key findings and contributions are summarized as follows: First, we show that many prominent LLMs struggle to capture the English color naming system, exhibiting a wide range of complexities that are often lower than the complexity of English. However, with increased size and instruction-tuning, LLMs can achieve high English-alignment and IB efficiency. Second, using our IICLL paradigm, we show that LLMs that perform well in the naming task are not merely mimicking patterns in their training data but are actually guided by a human-like inductive bias toward IB-efficiency. Specifically, we show that LLMs iteratively restructure initially random artificial systems towards greater IB-efficiency and increased human-alignment. However, among the models we tested, only the model with strongest in-context capabilities (Gemini 2.0) is able to recapitulate the wide range of near-optimal IB-tradeoffs observed in humans, while other models converge to low-complexity solutions. Finally, we show that Gemini can also develop structured category systems via IICLL in a domain that is qualitatively different from color, suggesting that our result could potentially apply also in other domains. Taken together, these findings demonstrate how human-aligned semantic categories can emerge in LLMs via the same fundamental principle that underlies semantic efficiency in humans. Importantly, neither humans nor LLMs are explicitly trained for optimizing the IB objective, suggesting that IB-efficiency may emerge to support intelligent behavior.
@@ 2 Background and motivation
Figure 1: (a) The standard WCS color naming grid (Kay2009World). (b) Color naming task with humans and LLMs. Multi-modal LLMs can observe colors either via text or images. (c) Illustration of the IICLL paradigm. At each generation $t$, an LLM is prompted with a small dataset for ICL, $d_{t-1}$, consisting of pairs of colors and pseudo labels sampled from the previous generation’s language, $L_{t-1}$. With these data in context, the LLM performs the naming task for the full space (a).
2.1 Why talk about color?
@@
For decades, cognitive scientists have used color as an essential tool to study perception and categorization. This is in part due to the unprecedented amount of human behavioral data available for color, which includes research on perceptual space, cross-linguistic semantic variation, category learning, and cultural evolution. Especially relevant to our study is the World Color Survey (WCS) dataset (Cook2005World), which contains color-naming data from 110 non-industrialized languages. Another relevant study by Xu2013Cultural demonstrated that the cultural transmission of initially random, artificial color-term systems in humans leads to more regular systems that resemble those documented in the WCS dataset. To our knowledge, these two data resources— on actual cross-linguistic naming patterns and on the cultural evolution of category systems—are unique to the domain of color, making it an ideal domain for evaluating how well LLMs align with human behavior. Furthermore, studying color naming in LLMs also has practical implications. Generative AI models, used for tasks like image generation and online product searches, require grounded representations of color language. In order for such models to interact in ways we expect them to, it is crucial to determine the extent to which state-of-the-art LLMs have learned the meaning of color terms that actually align with human naming patterns.
@@
While previous work has shown that human-aligned color representations can be recovered from LLMs (e.g., Abdou2021Can; Patel2022Mappinga), there is limited research on model behavior in the context of real-world settings (namely, prompt-based interactions). Particularly relevant to our work is a study by Marjieh2024Large that showed that several recent instruction-tuned models, namely GPT-3, its “ChatGPT” variants, GPT-3.5 and GPT-4 (Brown2020Language; OpenAI2024GPT4) and Mistral 7B Instruct (Jiang2023Mistral) can recover the English and Russian color naming systems by prompting the models to label hex codes. In contrast, we test a large set of 39 models with varying sizes and training stages, we consider both textual RGB inputs and image inputs (for multi-modal models), we analyze the LLMs’ color naming systems through the lens of the IB framework, and we further assess their underlying inductive learning biases in a cultural evolution process via iterated learning. We provide relevant background on the IB and iterated learning frameworks below.
2.2 The Information Bottleneck and semantic systems
@@
The IB framework for semantic systems (Zaslavsky2018Efficient), which we employ in this work, is based on the following communication model:a speaker wishes to communicate a mental representation, or belief state, $m\in\mathcal{M}$, defined as a probability distribution over world states $u\in\mathcal{U}$, by mapping it to a word $w\in\mathcal{W}$ via a stochastic encoder $q(w|m)$. A listener then receives $w$ and interprets the speaker’s intended meaning by constructing an estimator $\hat{m}_{w}$. In the case of color, for example, the world states $\mathcal{U}$ are given by a set of target colors (Figure 1a). To account for perceptual noise, it is assumed that each $m$ is a Gaussian distribution over the perceptual CIELAB color space centered around a corresponding target color (Zaslavsky2018Efficient). That is, each color is mentally represented by the speaker as a Gaussian distribution and the listener’s goal is to reconstruct the speaker’s belief state over colors.
@@
According to this framework, in order to communicate efficiently, the speaker and listener must jointly optimize the IB tradeoff between minimizing the complexity of their lexicon and maximizing its accuracy. Assuming the listener’s inferences are adapted to the speaker, the lexicon is defined by an encoder $q(w|m)$. Complexity in IB roughly corresponds to the number of bits required for communication, formally defined by $I_{q}(M;W)$, the mutual information between speaker’s meanings and words. Accuracy is defined by $I_{q}(W;U)$, the information that the speaker’s words maintain about the target world state, which is also inversely related to the KL-divergence between the speaker’s mental state and the listener’s inferred state, $\mathbb{E}_{q}[D[M\|\hat{M}$]] (Harremoes2007Information; Zaslavsky2020InformationTheoretic). An optimal lexicon, or semantic system, is one that minimizes the IB objective function
@@
This framework generates precise quantitative predictions that have been gaining converging empirical support across hundreds of languages and multiple semantic domains, ranging from perceptually-grounded domains such as color (Zaslavsky2018Efficient; Zaslavsky2022Evolution) to higher-level conceptual domains such as household objects (Zaslavsky2019Semantic; Taliaferro2025Bilingualsa) and personal pronouns (Zaslavsky2021Lets). In addition, it has been successfully applied to studying emergent communication in artificial agents (Chaabouni2022Emergent; Tucker2022Trading; Gualdoni2024Bridginga; Tucker2025HumanLike).
@@
As part of our evaluation of LLM color naming, we use the previously published IB color naming model from Zaslavsky2018Efficient. The black curve in Figure 3 shows the IB bound for color naming from this study, together with a reproduction of their results showing that color naming systems across languages (from the WCS as well as English from Lindsey2014Color) achieve near-optimal IB tradeoffs.
@@ 2.3 Iterated language learning
The second main theoretical framework we apply in this work is based on iterated learning (IL). IL is a paradigm in cognitive science for simulating cultural transmission and eliciting prior inductive biases (Griffiths2007Language; Kirby2008Cumulative; Griffiths2008Using). In a typical IL experiment, participants form chains of “generations.” At each generation $t$, a participant is exposed to data $d_{t-1}$ from the previous generation and then produces responses that become input for the next generation. In iterated language learning (ILL), participants learn examples of pairs of stimuli and artificial labels during a training period, after which they assign labels to new, unlabeled stimuli in the same meaning space. In doing so, participants produce a full category system $L_{t}$, which is sampled from to provide training examples to the next generation, and the process repeats. This process, which requires generalization from limited data, reveals learners’ inductive biases for certain linguistic or category structures (Griffiths2007Language; Griffiths2008Using). As shown in Griffiths2007Language, under certain conditions, namely that the IL agents are Bayesian who share priors and likelihood functions, this Markov chain converges to a stationary distribution over languages equal to the learners’ prior distribution $p(L)$. This makes the strong prediction that languages emerging from IL reflect the population’s underlying inductive biases. Although this is an asymptotic characterization of IL dynamics, in behavioral experiments with people, researchers observe rapid convergence to heavily biased distributions of languages (Kirby2015Compression).
@@
Previous research in machine learning has investigated dynamics related to ILL. In agent-based simulations, Ren2020Compositionala introduced the neural iterated learning (NIL) framework, and showed how it can lead to compositional language in neural network agents, and Carlsson2024Cultural showed that introducing communication-based training in NIL leads to IB-efficient color naming systems. These studies, however, did not explore LLMs as we do here. Zhu2024Eliciting adapted IL to LLMs with strong in-context learning capacities as a prompt-based workflow known as Iterated In-Context Learning (I-ICL) to elicit LLMs’ implicit prior distributions over aspects of world knowledge. I-ICL has recently been used to analyze cultural evolution in LLMs (Ren2024Bias) and to compare visual and linguistic abstractions between LLMs and humans (Kumar2024Comparinga). Here, we introduce iterated in-context language learning (IICLL, Figure 1c), which goes beyond prior work in leveraging the strong in-context learning abilities of LLMs to replicate as closely as possible the experimental conditions of ILL studies with humans, enabling a direct comparison to LLMs of their respective inductive biases, particularly with regards to semantic efficiency and alignment.
@@
The empirical comparison for our IICLL color naming study is the IL data from Xu2013Cultural. In their experiment, participants were asked to learn and transmit novel systems of color terms across thirteen generations. We focus on their main results that include twenty iterated learning chains, each initialized with a random partition of the WCS grid. These chains vary in the number of allowed color terms, ranging from two to six, and four replications of each condition. Participants were shown a set of randomly selected colors generated uniquely for each chain, and paired with corresponding pseudo words. After training, participants were asked to label all 330 colors of the WCS grid. Xu2013Cultural found that over time, the IL chains become increasingly regular and resembling the color naming systems documented in the WCS dataset. More recently, Imel2025Iterated found that these chains also converge to highly efficient systems along the IB bound. Figure 3 shows our reproduction of this finding (plotting only final generations).
@@ 3 Experimental setup
We consider 39 models across 6 model families: Gemini (Google2025Gemini), Gemma 3 (GemmaTeam2025Gemma), Llama 3 (Grattafiori2024Llama), Qwen 2.5 (Qwen2025Qwen25), Olmo 2 (OLMoTeam20252) and GPT-2 (Radford2019Languagea). Within each family, we vary models along several dimensions. Specifically, to gain insight into the properties that may influence the models’ behavior in our tasks, we consider models with different sizes, instruction-tuned versus base models, and text-based versus multi-modal models. For Olmo, we also considered its learning dynamics by analyzing training checkpoints. For a full list of models, see Table 1 in Appendix D.
@@
Following Zaslavsky2018Efficient, we use two main evaluation measures in our studies. First, the efficiency loss of a semantic system is measured by its minimum deviation from optimality, defined as $\varepsilon=\min_{\beta}\{\frac{1}{\beta}(\mathcal{F}_{\beta}[q]-\mathcal{F}_{\beta}^{*})\}$ where $\mathcal{F}_{\beta}^{*}$ is the optimal value of $\mathcal{F}_{\beta}$ in Eq. 1. Second, we measure the semantic (mis)alignment between two systems by the Normalized Information Distance (NID) (Kraskov2005Hierarchical; Vinh2010Information). NID is a metric capturing the distance between two clusterings (in this case, induced by color naming categories), providing a quantitative measure of structural similarity between the naming systems. Since NID is bounded in $[0,1]$, we take 1 - NID as a measure of similarity, or alignment. IB-alignment measures the similarity between a system and the nearest ($\epsilon$-fitted) optimal IB system. WCS-alignment measures the average alignment between a system and the WCS languages, and English-alignment measures the alignment between a system and English.

### v4

@@ 1 Introduction
As large language models (LLMs) become increasingly popular in everyday use, it is crucial to understand how their learning biases and representational capacities align with our own. Here, we investigate this by focusing on a key aspect of human intelligence: the ability to organize information into semantic categories (Croft, 2002; Rosch, 2002; Koch, 2008; Boster, 2005; Majid, 2015; Malt and Majid, 2013). This phenomenon presents two major challenges for AI. First, systems of semantic categories (semantic systems, for short) exhibit both universal patterns and cross-language differences (Berlin and Kay, 1969; Berlin, 1992; Croft, 2002), which LLMs must navigate. Second, LLMs are not grounded in the rich physical and social environment that humans are (Rosch, 1975; Labov, 1973), and it is unclear how these differences affect their ability to learn human-aligned semantic categories. Therefore, in order to understand whether LLMs can efficiently communicate with people and adapt to changing environments and communicative needs in a human-like manner, it is crucial to study whether LLMs are capable of structuring meaning according to the same principles that guide humans.
@@
To address this open gap in our understanding of LLMs, we propose a novel theoretical and cognitively-motivated framework for studying semantic systems in LLMs. We build on the framework of Zaslavsky et al. (2018), which argues that languages efficiently compress meanings into words by optimizing the Information Bottleneck (IB) principle (Tishby et al., 1999), instantiated as a tradeoff between the informational complexity and communicative accuracy of the lexicon. This framework has broad empirical support across human languages (Zaslavsky et al., 2018; 2019; 2021; Mollica et al., 2021; Zaslavsky et al., 2022). Furthermore, Imel et al. (2025) recently showed that a drive for IB-efficiency may be present in the individual inductive biases of human learners. While LLMs are trained on vast amounts of human language data encoded as text, they are not trained with respect to the IB objective. This raises the question: are LLMs capable of developing IB-efficient human-aligned semantic systems?
@@
We address this open question with an in-depth analysis in the domain of color—a key test case for categorization theories in cognitive science with rarely available human data as well as practical implication for human-LLM interactions (see Section 2.1)—and replicate with LLMs two influential human behavioral experiments (Figure 1). First, we conduct an English color naming experiment (analogous to the human experiment of Lindsey and Brown, 2014), designed to assess the efficiency and human-alignment of the color naming systems of LLMs. Second, we conduct an iterated learning experiment of pseudo color-naming systems (analogous to the human experiment of Xu et al., 2013), designed to reveal implicit inductive learning biases of LLMs by simulating a process of cultural transmission (see Section 2.3). For the latter, we extend Zhu and Griffiths (2024)’s iterated in-context learning (I-ICL) paradigm to iterated in-context language learning (IICLL).
@@
Our key findings and contributions are summarized as follows: First, we show that many prominent LLMs struggle to capture the English color naming system, exhibiting a wide range of complexities that are often lower than the complexity of English. However, with increased size and instruction-tuning, LLMs can achieve high English-alignment and IB efficiency. Second, using our IICLL paradigm, we show that LLMs that perform well in the naming task are not merely mimicking patterns in their training data but are actually guided by a human-like inductive bias toward IB-efficiency. Specifically, we show that LLMs iteratively restructure initially random artificial systems towards greater IB-efficiency and increased human-alignment. However, among the models we tested, only the model with strongest in-context capabilities (Gemini 2.0) is able to recapitulate the wide range of near-optimal IB-tradeoffs observed in humans, while other models converge to low-complexity solutions. Finally, we show that Gemini can also develop structured category systems via IICLL in a domain that is qualitatively different from color, suggesting that our result could potentially apply also in other domains.
Taken together, our findings demonstrate how human-aligned semantic categories can emerge in LLMs via the same fundamental principle that underlies semantic efficiency in humans. Importantly, neither humans nor LLMs are explicitly trained for optimizing the IB objective, suggesting that IB-efficiency, and optimal compression more generally, may emerge to support intelligent behavior.
@@ 2 Background and motivation
Figure 1: (a) The standard WCS color naming grid (Kay et al., 2009). (b) Color naming task with humans and LLMs. Multi-modal LLMs can observe colors either via text or images. (c) Illustration of the IICLL paradigm. At each generation $t$, an LLM is prompted with a small dataset for ICL, $d_{t-1}$, consisting of pairs of colors and pseudo labels sampled from the previous generation’s language, $L_{t-1}$. With these data in context, the LLM performs the naming task for the full space (a).
2.1 Talking about color
@@
For decades, cognitive scientists have used color as an essential tool to study perception and categorization. This is in part due to the unprecedented amount of human behavioral data available for color, which includes research on perceptual space, cross-linguistic semantic variation, category learning, and cultural evolution. Especially relevant to our study is the World Color Survey (WCS) dataset (Cook et al., 2005), which contains color-naming data from 110 non-industrialized languages. Another relevant study by Xu et al. (2013) demonstrated that the cultural transmission of initially random, artificial color-term systems in humans leads to more regular systems that resemble those documented in the WCS dataset. To our knowledge, these two data resources— on actual cross-linguistic naming patterns and on the cultural evolution of category systems—are unique to the domain of color, making it an ideal domain for evaluating how well LLMs align with human behavior. Furthermore, studying color naming in LLMs also has practical implications. Generative AI models, used for tasks like image generation and online product searches, require grounded representations of color language. In order for such models to interact in ways we expect them to, it is crucial to determine the extent to which state-of-the-art LLMs have learned the meaning of color terms that actually align with human naming patterns.
@@
While previous work has shown that human-aligned color representations can be recovered from LLMs (e.g., Abdou et al., 2021; Patel and Pavlick, 2022), there is limited research on model behavior in the context of real-world settings (namely, prompt-based interactions). Marjieh et al. (2024) showed that several recent instruction-tuned models (specifically, GPT-3, its “ChatGPT” variants, GPT-3.5 and GPT-4 (Brown et al., 2020; OpenAI et al., 2024) and Mistral 7B Instruct (Jiang et al., 2023)) can recover the English and Russian color naming systems by prompting the models to label hex codes. In contrast, we test a large set of 39 models with varying sizes and training stages, we consider both textual inputs and image inputs (for multi-modal models), we analyze the LLMs’ color naming systems through the lens of the IB framework, and we further assess their underlying inductive learning biases in a cultural evolution process via iterated learning. Next, we provide relevant background on the IB and iterated learning frameworks below.
2.2 The Information Bottleneck principle and semantic systems
@@
The IB framework for semantic systems (Zaslavsky et al., 2018), which we employ in this work, is based on the following communication model: a speaker wishes to communicate a mental representation, or belief state, $m\in\mathcal{M}$, defined as a probability distribution over world states $u\in\mathcal{U}$, by mapping it to a word $w\in\mathcal{W}$ via a stochastic encoder $q(w|m)$. A listener then receives $w$ and interprets the speaker’s intended meaning by constructing an estimator $\hat{m}_{w}$. In the case of color, for example, the world states $\mathcal{U}$ are given by a set of target colors (Figure 1a). To account for perceptual noise, it is assumed that each $m$ is a Gaussian distribution over the perceptual CIELAB color space centered around a corresponding target color (Zaslavsky et al., 2018). That is, each color is mentally represented by the speaker as a Gaussian distribution and the listener’s goal is to reconstruct the speaker’s belief state over colors.
@@
According to this framework, in order to communicate efficiently, the speaker and listener must jointly optimize the IB tradeoff between minimizing the complexity of their lexicon and maximizing its accuracy. Assuming the listener’s inferences are adapted to the speaker, the lexicon is defined by an encoder $q(w|m)$. Complexity in IB roughly corresponds to the number of bits required for communication, formally defined by $I_{q}(M;W)$, the mutual information between speaker’s meanings and words. Accuracy is defined by $I_{q}(W;U)$, the information that the speaker’s words maintain about the target world state, which is also inversely related to the KL-divergence between the speaker’s mental state and the listener’s inferred state, $\mathbb{E}_{q}[D[M\|\hat{M}$]] (Harremoes and Tishby, 2007; Zaslavsky, 2020). An optimal lexicon, or semantic system, is one that minimizes the IB objective function
@@
This framework generates precise quantitative predictions that have been gaining converging empirical support across hundreds of languages and multiple semantic domains, ranging from perceptually-grounded domains such as color (Zaslavsky et al., 2018; 2022) to higher-level conceptual domains such as household objects (Zaslavsky et al., 2019; Taliaferro et al., 2025) and personal pronouns (Zaslavsky et al., 2021). In addition, it has been successfully applied to studying emergent communication in artificial agents (Chaabouni et al., 2022; Tucker et al., 2022; Gualdoni et al., 2024; Tucker et al., 2025).
@@
As part of our evaluation of LLM color naming, we use the previously published IB color naming model from Zaslavsky et al. (2018). The black curve in Figure 3 shows the IB bound for color naming from this study, together with a reproduction of their results showing that color naming systems across languages (from the WCS as well as English from Lindsey and Brown (2014)) achieve near-optimal IB tradeoffs.
@@ 2.3 Iterated language learning
The second main theoretical framework we apply in this work is based on iterated learning (IL). IL is a paradigm in cognitive science for simulating cultural transmission and eliciting prior inductive biases (Griffiths and Kalish, 2007; Kirby et al., 2008; Griffiths et al., 2008). In a typical IL experiment, participants form chains of “generations.” At each generation $t$, a participant is exposed to data $d_{t-1}$ from the previous generation and then produces responses that become input for the next generation. In iterated language learning (ILL), participants learn examples of pairs of stimuli and artificial labels during a training period, after which they assign labels to new, unlabeled stimuli in the same meaning space. In doing so, participants produce a full category system $L_{t}$, which is sampled from to provide training examples to the next generation, and the process repeats. This process, which requires generalization from limited data, reveals learners’ inductive biases for certain linguistic or category structures (Griffiths and Kalish, 2007; Griffiths et al., 2008). As shown in Griffiths and Kalish (2007), under certain conditions, namely that the IL agents are Bayesian who share priors and likelihood functions, this Markov chain converges to a stationary distribution over languages equal to the learners’ prior distribution $p(L)$. This makes the strong prediction that languages emerging from IL reflect the population’s underlying inductive biases. Although this is an asymptotic characterization of IL dynamics, in behavioral experiments with people, researchers observe rapid convergence to highly non-uniform distributions of languages (Kirby et al., 2015).
@@
Previous research in machine learning has investigated dynamics related to ILL. In agent-based simulations, Ren et al. (2020) introduced the neural iterated learning (NIL) framework, and showed how it can lead to compositional language in neural network agents, and Carlsson et al. (2024) showed that introducing communication-based training in NIL leads to IB-efficient color naming systems. These studies, however, did not explore LLMs as we do here. Zhu and Griffiths (2024) adapted IL to LLMs with strong in-context learning capacities as a prompt-based workflow known as Iterated In-Context Learning (I-ICL) to elicit LLMs’ implicit prior distributions over aspects of world knowledge. I-ICL has recently been used to analyze cultural evolution in LLMs (Ren et al., 2024) and to compare visual and linguistic abstractions between LLMs and humans (Kumar et al., 2024). Here, we introduce iterated in-context language learning (IICLL, Figure 1c), which goes beyond prior work in leveraging the strong in-context learning abilities of LLMs to replicate as closely as possible the experimental conditions of ILL studies with humans, enabling a direct comparison to LLMs of their respective inductive biases, particularly with regards to semantic efficiency and alignment.
@@
The empirical comparison for our IICLL color naming study is the IL data from Xu et al. (2013). In their experiment, participants were asked to learn and transmit novel systems of color terms across thirteen generations. We focus on their main results that include twenty iterated learning chains, each initialized with a random partition of the WCS grid. These chains vary in the number of allowed color terms, ranging from two to six, and four replications of each condition. Participants were shown a set of randomly selected colors generated uniquely for each chain, and paired with corresponding pseudo words. After training, participants were asked to label all 330 colors of the WCS grid. Xu et al. (2013) found that over time, the IL chains become increasingly regular and resemble the color naming systems documented in the WCS dataset. More recently, Imel et al. (2025) found that these chains also converge to highly efficient systems along the IB bound. Figure 3 shows our reproduction of this finding (plotting only final generations).
@@ 3 Experimental setup
We consider 39 models across 6 model families: Gemini (Google, 2025), Gemma 3 (Gemma-Team et al., 2025), Llama 3 (Grattafiori et al., 2024), Qwen 2.5 (Qwen et al., 2025), Olmo 2 (OLMo-Team et al., 2025) and GPT-2 (Radford et al., 2019). Within each family, we vary models along several dimensions. Specifically, to gain insight into the properties that may influence the models’ behavior in our tasks, we consider models with different sizes, instruction-tuned versus base models, and text-based versus multi-modal models. For Olmo, we also considered its learning dynamics by analyzing training checkpoints. For a full list of models, see Table 1 in Appendix D.
@@
Following Zaslavsky et al. (2018), we use two main evaluation measures in our studies. First, the efficiency loss of a semantic system is measured by its minimum deviation from optimality, defined as $\varepsilon=\min_{\beta}\{\frac{1}{\beta}(\mathcal{F}_{\beta}[q]-\mathcal{F}_{\beta}^{*})\}$ where $\mathcal{F}_{\beta}^{*}$ is the optimal value of $\mathcal{F}_{\beta}$ in Eq. 1. Second, we measure the semantic (mis)alignment between two systems by the Normalized Information Distance (NID) (Kraskov et al., 2005; Vinh et al., 2010). NID is a metric capturing the distance between two clusterings (in this case, induced by color naming categories), providing a quantitative measure of structural similarity between the naming systems. Since NID is bounded in $[0,1]$, we take 1 - NID as a measure of similarity, or alignment. IB-alignment measures the similarity between a system and the nearest ($\epsilon$-fitted) optimal IB system. WCS-alignment measures the average alignment between a system and the WCS languages, and English-alignment measures the alignment between a system and English.